# Scientific Machine Learning of Chaotic Systems Discovers Governing Equations for Neural Populations

### v2

@@ III Discussion
Here, we introduce a new method to fit universal differential equations in chaotic systems and demonstrate that it learns novel governing equations for neural population activity. The approach harnesses the prediction error method to train the UDE system, from which we then extract a symbolic form of the learned dynamics via symbolic regression [10]. We demonstrate that our method can learn the dynamics of a classical chaotic system, using the Rössler system [22]. A standard UDE alone cannot learn this system effectively. However, our combined PEM-UDE approach captures the dynamics accurately, demonstrating that our method has identified the true functional form of the dynamics rather than merely making projections within conventional statistical boundaries. Prior work in UDEs has focused on stiff systems [7], fitting UDEs in the presence of stochastic noise [69], and extracting explainable functions from UDE networks [16]. Independently, prior work in parameter optimization using the PEM has shown its utility in fitting parameters of dynamical systems [17, 52], and has shown that it can even outperform neural network fits of dynamics [70, 71]. Our approach here represents the first time the PEM has been used to train UDEs and to extract a functional form of these UDEs to provide an analytical result. Our approach succeeds where standard UDE methods fail because it effectively smooths the parameter landscape the UDE must optimize. Importantly, this smoothing removes a sensitive dependence on the parameters of the system while retaining a loss minimum at the same value as the original system, creating a well-behaved problem when fitting the dynamics. We also extend this application to a second chaotic system from an electrical circuit [72, 53], and show that even in the presence of significant observational noise, the PEM-UDE method can recover the true functional form of an unknown state. A more accurate estimation of the parameters of these recovered dynamics could be achieved by taking the functional form learned from the PEM-UDE dynamics and applying a more advanced optimization technique that fully uses the ability to simulate the system in different regimes (e.g., simulation-based inference [55]).
@@
Having shown that the PEM-UDE approach can learn the dynamics of chaotic systems, we then consider how to learn governing equations in systems of neurons under biologically plausible constraints. These constraints are of two varieties: experimental constraints on the data available, including using only data sources that can be feasibly acquired in vivo (firing rate, membrane voltage); and neurobiological, where the sparsity of interneuron connections can be accurately modeled. Experimental constraints are useful to bear in mind as they can make the method more immediately applicable, but advances in methods can often relax these constraints over time. Biological constraints, and connection density in particular, are more important to accurately model, as they fulfill an important role in the types of information processing possible within the brain [49, 50]. We use a missing-physics approach, where the initial form of the population activity is given by a next-generation neural mass model [42, 43]. NGNMMs have shown great promise in capturing emergent properties of neuron populations while retaining the microscale parameters that are typically manipulated in vitro [44, 73], but rely on the assumption of all-to-all connectivity that is inherently non-physiological for most brain regions [42]. There have been varied attempts to sidestep this limitation and include corrections for sparsity by reweighting the degree of synaptic conductance or the distribution of synaptic currents, with varying degrees of success [47, 59, 74, 60]. Each of these approaches comes with new assumptions and its own limitations. Here, we avoid this by learning a novel NGNMM as the governing equation for a population of Izhikevich neurons directly from the observations of the network itself. We demonstrate that the governing equations learned by this approach predict novel emergent properties that are observed directly in biology, like the shift from $\theta$ to $\alpha$ as the dominant frequency band as networks become more sparse – a feature seen in the brain moving from densely connected regions (e.g., hippocampus) to sparsely connected regions (e.g., prefrontal cortex) [61, 62]. These governing equations also predict that more densely connected regions will exhibit greater intraregional synchrony at rest than sparsely connected networks. To test this hypothesis, we compute the intraregional synchrony from three datasets of intracranial electrode recordings (rat MEC, rat prefrontal cortex, and human orbitofrontal cortex) [63, 65, 67]. Across all datasets, we see that the pattern predicted by the discovered equations holds true: intraregional synchrony at rest is inversely correlated with the degree of connectivity between neurons in the region, giving us further validation that the PEM-UDE method is capable of uncovering equations that accurately describe complex biological systems.
@@
Despite its strengths, the method presented here also has limitations that will be addressed in future work. Extremely noisy systems can be difficult to learn, especially as chaotic systems can be complicated to average over trials. Additionally, care must be taken because it is possible that nonphysical solutions also shadow physical ones [24]. The inclusion of information during training via the PEM and the extraction of a symbolic form of the dynamics help mitigate these difficulties. In models that include known physics, PEM can lead to close-but-inaccurate minima if the parameters are not exactly known. This limits the method’s applicability to scenarios where the parameters included in the known physical model are sufficiently accurate. In principle, this can be overcome by also fitting to a range of simulations across these parameters, but the training time will be significantly increased. Finally, the method presented here is best suited for situations with a known structure of data (i.e., some driving inputs are known). Neural experiments frequently provide this level of accuracy, but there is also substantial interest in resting state data [75]. Adapting our approach to accurately fit stationary time series without known perturbations is thus also an important extension for neuroscience applications.
@@
Although the work presented here demonstrates the significant utility of this approach in neuroscience, there are notable areas for further development. Future work can incorporate loss objectives that include optimization of initial conditions, as this can assist in generalization to real-world systems [70, 76]. Future work can also focus on learning ”black-box” representations of the dynamics, where symbolic regression methods are not applied to learn the dynamics. This can be useful in situations where complexity penalties (i.e., enforced sparsity within the learned expression) bias away from the correct dynamics. For example, in a case where many small inputs lead to complex interactions based on their temporal pattern, the high-dimensional representation present in the neural network may be more desirable. The temporal sequences of visual cortex projections to the prefrontal cortex are an example of this kind of structured high-dimensional input space [48]. In such a case, rather than fitting the dynamics symbolically, a second optimization run of the UDE without PEM after the initial combined run could lead to more generalizable results (as this forces the contribution of PEM to 0 rather than just naturally minimizing it over the optimization problem). Including a method for directly estimating stochastic processes within the dynamics, as has been done in some standard UDE approaches [69], would provide an additional layer of generalization to this approach. Finally, extending the PEM-UDE approach to systems of partial differential equations would allow for a direct fitting of neural field models under differing degrees of connection density, which are of increasing importance as neuroscience begins to understand the importance of traveling waves in the brain [77, 19] and how the interplay between densely and sparsely connected regions structures the flow of cognitive processes [49].
@@ IV.1 Scientific machine learning in the Julia ecosystem
The Julia ecosystem has a robust infrastructure for scientific machine learning [16] that we use to build the PEM-UDE method described in this work. As setting up and training UDEs in Julia has been discussed in other publications ([16, 7] are notable examples), we encourage interested readers to simply peruse the code to see the details of the implementation and content ourselves with a few brief technical notes here on details about the choices of solvers, network structure, fitting procedure, and symbolic regression.
@@
We choose the radial basis function (RBF; notation given in [78]) as the activation function for these layers, given by
@@
The RBF is selected for both theoretical and practical reasons. There is a robust body of literature showing that the RBF is useful in the context of universal approximation [78, 79, 80]. Because a neural network of this form can approximate any function given a sufficient depth of the network [78], the choice of network size is therefore arbitrary beyond a specific size. It can be set freely based on the availability of computational resources for training. For the networks shown in this work, we choose a depth of four layers, where for an input dimension $N_{1}$ and output dimension $N_{2}$, each network is of size $N_{1}\times 10\times 10\times N_{2}$.
@@
As the PEM portion of training requires interpolating the original training data to the solver time step, we use DataInterpolations.jl to do simple linear interpolation between observations [81]. Although there are many optimizers provided to train the UDEs, we have defaulted to using the AdamW method (decoupled weight decay and gradient updates; [82]) in all of the results presented here.
@@
While the dynamics fit by the combined PEM-UDE method appear to capture the true dynamics of the system, it is important to note that they are trained on limited data, and evaluating robustness – particularly in chaotic systems with their sensitivity to initial conditions – is challenging to carry out. The approach to mitigating this is to fit a symbolic expression to the dynamics of the UDE networks that we train to provide a form of the dynamics that can be interrogated with standard analytical approaches. For the purposes of this paper, we do not make any claims as to the best way of extracting the symbolic form of the UDE network and note that there are advantages to different approaches [3]. We provide examples of two of these approaches [10, 14] in the code repository associated with this work. As the STLSQ approach is significantly faster than the genetic algorithm approach when large amounts of data ($N\gtrapprox 10,000$ data points) are required to fit the functional form accurately [14], we default to this method for the results shown here. In principle, however, the trained UDE can be used to generate a subsample of arbitrary density, so either approach can be used, depending on the requirements of the problem at hand.
@@
The notation for the symbolic recovery of terms from neural networks is adapted from [10], so we briefly explain it here. $\Theta$ is defined as the feature library with some linear combination of features given by weights $\xi$ so that $y=\Theta\xi$. To reduce the library components that constitute the functional form, a sparse solution is found via LASSO [83]. The objective function is therefore given by
$\operatorname*{arg\,min}_{\xi}\frac{1}{2}||\Theta\xi-y||_{2}+\lambda||\xi||_{1}$ (10)
@@
In the implementation used in this work, this objective is approximated by sequential thresholded least-squares (STLSQ) solved for using the alternating direction method of multipliers (ADMM; [84]), as implemented in DataDrivenDiffEq.jl [56].
@@ IV.2 Background on the Rössler system
For the results presented in the main text, we select $K=0.25$. This choice is driven by the steepness of the parameter landscapes as discussed in the results, and other reasonable values of this hyperparameter will not affect the accuracy of the learned dynamics, provided an appropriate solver is chosen. The results presented here are based on 100,000 training epochs with the AdamW optimizer [82]. The loss function used to train the UDE network is simply mean-squared error between the predicted states at each time step and the true time series (with a small amount of noise added for robustness during training):
$\ell(\theta)=||\hat{\mathbf{x}}-\mathbf{x}||^{2}$ (13)
For the parameters selected in our analyses, we characterized the degree of chaos via the Lyapunov time of the system as $\lambda^{-1}\approx 10$. Most parameter combinations in the surrounding space have similar values ($>80\%$ have $\lambda^{-1}\in[0,25]$ see supplemental code for computation of the empirical CDF of the maximum Lyapunov exponents). Lyapunov exponents were computed using standard functions offered in ChaosTools.jl [85, 86].
@@ IV.3 Background on Petrzela-Polak circuit
Petrzela and Polak [72] introduced a category of analog electrical circuits exhibiting chaotic dynamics under three constraints: a canonical form, structurally stable attractors, and novel dynamics. Here, we work with a particular realization that provides chaotic dynamics that are visually interesting and can be physically realized in a simple circuit, with the attractor seen on an oscilloscope (demonstrated in [53], from which we take the parameters and non-dimensional form discussed here). For the circuit shown in 3a, the equations describing the three observed states are given by
@@
All parameters for both the dimensional and dimensionless systems are taken from [53], and are provided in Extended Data Table 2.
@@ IV.4 Spiking neural networks and next-generation neural mass models
Next-generation neural mass models rely on the Ott-Antonsen (OA) ansatz to derive the mean-field activity of a population of neurons [45, 42]. Briefly, this approach was introduced as a means of bypassing the need to set up an entire Fokker-Planck equation to derive the mean-field activity of a population of coupled oscillators, provided certain key assumptions are met. The OA ansatz [45] relies on the assumption that the population of coupled oscillators being examined exists in the thermodynamic limit (i.e., $N\rightarrow\infty$) and that the probability distribution of the observable (typically voltage) in phase space converges to a Lorentzian. However, this derivation can approximate other distributions to a reasonable degree of accuracy [47]. The OA ansatz approach also assumes that the noise input into each neuron receives a heterogeneous, quenched noise input $\eta$ drawn from a general distribution. The key point of this assumption is that the currents cannot be stochastic; otherwise, the prior assumption (the Lorentzian ansatz) is violated, which is further discussed below. The OA ansatz also assumes that the conditional probabilities of the macroscopic variables are separable. This assumption and its limitations have been discussed at length [74], and a second-order circular cumulant approximation has been derived for neural mass models [59]. Finally, the OA ansatz makes the assumption we address in this paper: The network being approximated has an all-to-all connectivity structure. As we have discussed above, this is a fundamentally non-physiological assumption for most regions of the brain, and removing this assumption is an important step in developing more accurate neural mass models.
As we do not require the OA ansatz for the work presented here beyond the initial assumption of the functional form for population activity that will be used to train the UDE, we omit most of the derivation and discuss only the key points. The equations shown here are adapted from earlier work that shows the entire derivation [43]. Following this work, we simulate a network of Izhikevich neurons that take the form
@@
$I_{\text{syn,}j}$ $=g_{\text{syn}}s(e_{r}-V_{j})$ (19)
$\dot{s_{j}}$ $=-\frac{s_{j}}{\tau_{s}}+\frac{s_{\text{jump}}}{N}\sum_{j=1}^{N}{\sum_{t_{j}^{k}<t}\delta(t-t_{j}^{k})}$ (20)
@@
where $r$ is the population firing rate, $v$ is the mean voltage, $w$ the mean recovery current, and $s$ the mean synaptic activity. As discussed in earlier work, this is an accurate representation of the mean-field activity for fully connected networks [43].
For all the simulations presented here, we simulate a network of 1,000 Izhikevich neurons with varying degrees of network sparsity, where the probability of two neurons forming a synapse is given by $p(c)\in[0.05,0.1,0.2,...1.0]$ as indicated for the sparsities shown in Fig. 4. In a standard network of neurons, as the sparsity of connections is varied, it is necessary to rescale the synaptic currents proportionally with the median in-degree of connectivity $K$ to maintain balanced (synchronous) activity. Previous work has shown that rescaling the synaptic conductance with $g\propto\frac{1}{\sqrt{K}}$ – with similar rescaling of background currents, if present – provides an accurate scaling for a broad spectrum of network arrangements (in-degrees with Lorentzian, Gaussian, and Erdős-Rényi distributions) [47]. Following this work, we rescale synaptic conductance with $p(c)$ such that
@@
As in the previous sections, we set up the UDE to learn the unknown terms for the states we are interested in, but with two key differences. We are learning terms for both the first and second states in this derivation and thus applying PEM to both of these states as well. We also modify Eq. 13 to train on only data for the first two states, as these are the only states typically observable in experimental data. The initial parameters for this system are given in Table 3, and the formal description of the PEM-UDE system is given by
@@
$\dot{r}$ $=\frac{\Delta}{\pi}+2rv-(\alpha+g_{\text{syn}}s)r+f_{1}(\overset{\rightharpoonup}{u},p_{c})$ (24a)
$\dot{v}$ $=v(v-\alpha)-w+\bar{\eta}+I_{\text{ext}}+g_{\text{syn}}s(e_{r}-v)-(\pi r)^{2}+f_{2}(\overset{\rightharpoonup}{u},p_{c})$ (24b)
@@
$f_{1}(\overset{\rightharpoonup}{u},p_{c})$ $=\phi_{1}s+\phi_{2}p_{c}+\phi_{3}vp_{c}+c_{1}$ (24e)
$f_{2}(\overset{\rightharpoonup}{u},p_{c})$ $=\phi_{4}r+\phi_{5}w^{2}+c_{2}$ (24f)
@@
$b$0.0062 Dimensionless Izhikevich recovery current parameter
@@
Table 3: Parameters of the Izhikevich neuron population and NGNMM, from [43].
@@
$\phi_{1}$0.45 Reduction of synaptic activity contribution to firing rate
$\phi_{2}$0.20 Reduction of mean firing rate as connectivity decreases
$\phi_{3}$ 0.55 Increased voltage contribution to firing rate as connectivity decreases
$\phi_{4}$ 0.89 Increased contribution of firing rate to voltage
$\phi_{5}$ 1.72 Starkly increased contribution of recovery current to voltage
$c_{1}$ 0.17 Rescaling of $\Delta$
$c_{2}$0.14 Rescaling of $\bar{\eta}$
Table 4: Parameters of the symbolic PEM-UDE learned NGNMM.

### v3

@@ III Discussion
Here, we introduce a new method to fit universal differential equations in chaotic systems and demonstrate that it learns novel governing equations for neural population activity. The approach harnesses the prediction error method to train the UDE system, from which we then extract a symbolic form of the learned dynamics via symbolic regression [brunton2016]. We demonstrate that our method can learn the dynamics of a classical chaotic system, using the Rössler system [rossler1976]. A standard UDE alone cannot learn this system effectively. However, our combined PEM-UDE approach captures the dynamics accurately, demonstrating that our method has identified the true functional form of the dynamics rather than merely making projections within conventional statistical boundaries. Prior work in UDEs has focused on stiff systems [kim2021], fitting UDEs in the presence of stochastic noise [oleary2022], and extracting explainable functions from UDE networks [rackauckas2021]. Independently, prior work in parameter optimization using the PEM has shown its utility in fitting parameters of dynamical systems [ljung2002, larsson2009], and has shown that it can even outperform neural network fits of dynamics [sun2021, pillonetto2025]. Our approach here represents the first time the PEM has been used to train UDEs and to extract a functional form of these UDEs to provide an analytical result. Our approach succeeds where standard UDE methods fail because it effectively smooths the parameter landscape the UDE must optimize. Importantly, this smoothing removes a sensitive dependence on the parameters of the system while retaining a loss minimum at the same value as the original system, creating a well-behaved problem when fitting the dynamics. We also extend this application to a second chaotic system from an electrical circuit [petrzela2019, sprott2022], and show that even in the presence of significant observational noise, the PEM-UDE method can recover the true functional form of an unknown state. A more accurate estimation of the parameters of these recovered dynamics could be achieved by taking the functional form learned from the PEM-UDE dynamics and applying a more advanced optimization technique that fully uses the ability to simulate the system in different regimes (e.g., simulation-based inference [goncalves2020]).
@@
Having shown that the PEM-UDE approach can learn the dynamics of chaotic systems, we then consider how to learn governing equations in systems of neurons under biologically plausible constraints. These constraints are of two varieties: experimental constraints on the data available, including using only data sources that can be feasibly acquired in vivo (firing rate, membrane voltage); and neurobiological, where the sparsity of interneuron connections can be accurately modeled. Experimental constraints are useful to bear in mind as they can make the method more immediately applicable, but advances in methods can often relax these constraints over time. Biological constraints, and connection density in particular, are more important to accurately model, as they fulfill an important role in the types of information processing possible within the brain [balcioglu2023, fruengel2025]. We use a missing-physics approach, where the initial form of the population activity is given by a next-generation neural mass model [montbrio2015, chen2022]. NGNMMs have shown great promise in capturing emergent properties of neuron populations while retaining the microscale parameters that are typically manipulated in vitro [coombes2019, sheheitli2024], but rely on the assumption of all-to-all connectivity that is inherently non-physiological for most brain regions [montbrio2015]. There have been varied attempts to sidestep this limitation and include corrections for sparsity by reweighting the degree of synaptic conductance or the distribution of synaptic currents, with varying degrees of success [divolo2018, goldobin2021, divolo2022, goldobin2019, clusella2024]. Each of these approaches comes with new assumptions and its own limitations. Here, we avoid this by learning a novel NGNMM as the governing equation for a population of Izhikevich neurons directly from the observations of the network itself. We demonstrate that the governing equations learned by this approach predict novel emergent properties that are observed directly in biology, like the shift from $\theta$ to $\alpha$ as the dominant frequency band as networks become more sparse – a feature seen in the brain moving from densely connected regions (e.g., hippocampus) to sparsely connected regions (e.g., prefrontal cortex) [lubenov2009, halgren2019]. These governing equations also predict that more densely connected regions will exhibit greater intraregional synchrony at rest than sparsely connected networks. To test this hypothesis, we compute the intraregional synchrony from three datasets of intracranial electrode recordings (rat MEC, rat prefrontal cortex, and human orbitofrontal cortex) [hernandez-perez2020, peyrache2009, saez2018]. Across all datasets, we see that the pattern predicted by the discovered equations holds true: intraregional synchrony at rest is inversely correlated with the degree of connectivity between neurons in the region, giving us further validation that the PEM-UDE method is capable of uncovering equations that accurately describe complex biological systems.
@@
Despite its strengths, the method presented here also has limitations that will be addressed in future work. Extremely noisy systems can be difficult to learn, especially as chaotic systems can be complicated to average over trials. Additionally, care must be taken because it is possible that nonphysical solutions also shadow physical ones [chandramoorthy2021]. The inclusion of information during training via the PEM and the extraction of a symbolic form of the dynamics help mitigate these difficulties. In models that include known physics, PEM can lead to close-but-inaccurate minima if the parameters are not exactly known. This limits the method’s applicability to scenarios where the parameters included in the known physical model are sufficiently accurate. In principle, this can be overcome by also fitting to a range of simulations across these parameters, but the training time will be significantly increased. Finally, the method presented here is best suited for situations with a known structure of data (i.e., some driving inputs are known). Neural experiments frequently provide this level of accuracy, but there is also substantial interest in resting state data [nozari2024]. Adapting our approach to accurately fit stationary time series without known perturbations is thus also an important extension for neuroscience applications.
@@
Although the work presented here demonstrates the significant utility of this approach in neuroscience, there are notable areas for further development. Future work can incorporate loss objectives that include optimization of initial conditions, as this can assist in generalization to real-world systems [sun2021, ji2021]. Future work can also focus on learning ”black-box” representations of the dynamics, where symbolic regression methods are not applied to learn the dynamics. This can be useful in situations where complexity penalties (i.e., enforced sparsity within the learned expression) bias away from the correct dynamics. For example, in a case where many small inputs lead to complex interactions based on their temporal pattern, the high-dimensional representation present in the neural network may be more desirable. The temporal sequences of visual cortex projections to the prefrontal cortex are an example of this kind of structured high-dimensional input space [pathak2024]. In such a case, rather than fitting the dynamics symbolically, a second optimization run of the UDE without PEM after the initial combined run could lead to more generalizable results (as this forces the contribution of PEM to 0 rather than just naturally minimizing it over the optimization problem). Including a method for directly estimating stochastic processes within the dynamics, as has been done in some standard UDE approaches [oleary2022], would provide an additional layer of generalization to this approach. Finally, extending the PEM-UDE approach to systems of partial differential equations would allow for a direct fitting of neural field models under differing degrees of connection density, which are of increasing importance as neuroscience begins to understand the importance of traveling waves in the brain [roberts2019, breakspear2017] and how the interplay between densely and sparsely connected regions structures the flow of cognitive processes [balcioglu2023].
@@ IV.1 Scientific machine learning in the Julia ecosystem
The Julia ecosystem has a robust infrastructure for scientific machine learning [rackauckas2021] that we use to build the PEM-UDE method described in this work. As setting up and training UDEs in Julia has been discussed in other publications ([rackauckas2021, kim2021] are notable examples), we encourage interested readers to simply peruse the code to see the details of the implementation and content ourselves with a few brief technical notes here on details about the choices of solvers, network structure, fitting procedure, and symbolic regression.
@@
We choose the radial basis function (RBF; notation given in [craddock1996]) as the activation function for these layers, given by
@@
The RBF is selected for both theoretical and practical reasons. There is a robust body of literature showing that the RBF is useful in the context of universal approximation [craddock1996, park1991, zhao2019]. Because a neural network of this form can approximate any function given a sufficient depth of the network [craddock1996], the choice of network size is therefore arbitrary beyond a specific size. It can be set freely based on the availability of computational resources for training. For the networks shown in this work, we choose a depth of four layers, where for an input dimension $N_{1}$ and output dimension $N_{2}$, each network is of size $N_{1}\times 10\times 10\times N_{2}$.
@@
As the PEM portion of training requires interpolating the original training data to the solver time step, we use DataInterpolations.jl to do simple linear interpolation between observations [bhagavan2024]. Although there are many optimizers provided to train the UDEs, we have defaulted to using the AdamW method (decoupled weight decay and gradient updates; [loshchilov2019]) in all of the results presented here.
@@
While the dynamics fit by the combined PEM-UDE method appear to capture the true dynamics of the system, it is important to note that they are trained on limited data, and evaluating robustness – particularly in chaotic systems with their sensitivity to initial conditions – is challenging to carry out. The approach to mitigating this is to fit a symbolic expression to the dynamics of the UDE networks that we train to provide a form of the dynamics that can be interrogated with standard analytical approaches. For the purposes of this paper, we do not make any claims as to the best way of extracting the symbolic form of the UDE network and note that there are advantages to different approaches [cranmer2020]. We provide examples of two of these approaches [brunton2016, cranmer2023] in the code repository associated with this work. As the STLSQ approach is significantly faster than the genetic algorithm approach when large amounts of data ($N\gtrapprox 10,000$ data points) are required to fit the functional form accurately [cranmer2023], we default to this method for the results shown here. In principle, however, the trained UDE can be used to generate a subsample of arbitrary density, so either approach can be used, depending on the requirements of the problem at hand.
@@
The notation for the symbolic recovery of terms from neural networks is adapted from [brunton2016], so we briefly explain it here. $\Theta$ is defined as the feature library with some linear combination of features given by weights $\xi$ so that $y=\Theta\xi$. To reduce the library components that constitute the functional form, a sparse solution is found via LASSO [tibshirani1996]. The objective function is therefore given by
[…]
@@
In the implementation used in this work, this objective is approximated by sequential thresholded least-squares (STLSQ) solved for using the alternating direction method of multipliers (ADMM; [boyd2010]), as implemented in DataDrivenDiffEq.jl [juliusmartensen2021].
@@ IV.2 Background on the Rössler system
For the results presented in the main text, we select $K=0.25$. This choice is driven by the steepness of the parameter landscapes as discussed in the results, and other reasonable values of this hyperparameter will not affect the accuracy of the learned dynamics, provided an appropriate solver is chosen. The results presented here are based on 100,000 training epochs with the AdamW optimizer [loshchilov2019]. The loss function used to train the UDE network is simply mean-squared error between the predicted states at each time step and the true time series (with a small amount of noise added for robustness during training):
[…]
For the parameters selected in our analyses, we characterized the degree of chaos via the Lyapunov time of the system as $\lambda^{-1}\approx 10$. Most parameter combinations in the surrounding space have similar values ($80\%$ have $\lambda^{-1}\in[0,25]$ - see supplemental code for computation of the empirical CDF of the maximum Lyapunov exponents). Lyapunov exponents were computed using standard functions offered in ChaosTools.jl [datseris2018, datseris2022].
@@ IV.3 Background on Petrzela-Polak circuit
Petrzela and Polak [petrzela2019] introduced a category of analog electrical circuits exhibiting chaotic dynamics under three constraints: a canonical form, structurally stable attractors, and novel dynamics. Here, we work with a particular realization that provides chaotic dynamics that are visually interesting and can be physically realized in a simple circuit, with the attractor seen on an oscilloscope (demonstrated in [sprott2022], from which we take the parameters and non-dimensional form discussed here). For the circuit shown in 3a, the equations describing the three observed states are given by
@@
All parameters for both the dimensional and dimensionless systems are taken from [sprott2022], and are provided in Extended Data Table 2.
@@ IV.4 Spiking neural networks and next-generation neural mass models
Next-generation neural mass models rely on the Ott-Antonsen (OA) ansatz to derive the mean-field activity of a population of neurons [ott2008a, montbrio2015]. Briefly, this approach was introduced as a means of bypassing the need to set up an entire Fokker-Planck equation to derive the mean-field activity of a population of coupled oscillators, provided certain key assumptions are met. The OA ansatz [ott2008a] relies on the assumption that the population of coupled oscillators being examined exists in the thermodynamic limit (i.e., $N\rightarrow\infty$) and that the probability distribution of the observable (typically voltage) in phase space converges to a Lorentzian. However, this derivation can approximate other distributions to a reasonable degree of accuracy [divolo2018]. The OA ansatz approach also assumes that the noise input into each neuron receives a heterogeneous, quenched noise input $\eta$ drawn from a general distribution. The key point of this assumption is that the currents cannot be stochastic; otherwise, the prior assumption (the Lorentzian ansatz) is violated, which is further discussed below. The OA ansatz also assumes that the conditional probabilities of the macroscopic variables are separable. This assumption and its limitations have been discussed at length [goldobin2019], and a second-order circular cumulant approximation has been derived for neural mass models [divolo2022]. Finally, the OA ansatz makes the assumption we address in this paper: The network being approximated has an all-to-all connectivity structure. As we have discussed above, this is a fundamentally non-physiological assumption for most regions of the brain, and removing this assumption is an important step in developing more accurate neural mass models.
As we do not require the OA ansatz for the work presented here beyond the initial assumption of the functional form for population activity that will be used to train the UDE, we omit most of the derivation and discuss only the key points. The equations shown here are adapted from earlier work that shows the entire derivation [chen2022]. Following this work, we simulate a network of Izhikevich neurons that take the form
@@
$I_{\text{syn,}j}$ $=g_{\text{syn}}s(e_{r}-V_{j})$ (19a)
$\dot{s_{j}}$ $=-\frac{s_{j}}{\tau_{s}}+\frac{s_{\text{jump}}}{N}\sum_{j=1}^{N}{\sum_{t_{j}^{k}<t}\delta(t-t_{j}^{k})}$ (19b)
@@
where $r$ is the population firing rate, $v$ is the mean voltage, $w$ the mean recovery current, and $s$ the mean synaptic activity. As discussed in earlier work, this is an accurate representation of the mean-field activity for fully connected networks [chen2022].
For all the simulations presented here, we simulate a network of 1,000 Izhikevich neurons with varying degrees of network sparsity, where the probability of two neurons forming a synapse is given by $p(c)\in[0.05,0.1,0.2,...1.0]$ as indicated for the sparsities shown in Fig. 4. In a standard network of neurons, as the sparsity of connections is varied, it is necessary to rescale the synaptic currents proportionally with the median in-degree of connectivity $K$ to maintain balanced (synchronous) activity. Previous work has shown that rescaling the synaptic conductance with $g\propto\frac{1}{\sqrt{K}}$ – with similar rescaling of background currents, if present – provides an accurate scaling for a broad spectrum of network arrangements (in-degrees with Lorentzian, Gaussian, and Erdős-Rényi distributions) [divolo2018]. Following this work, we rescale synaptic conductance with $p(c)$ such that
@@
with $s_{\text{jump},0}=s_{\text{jump}}\sqrt{1.5}\approx 1.5$. As in the previous sections, we set up the UDE to learn the unknown terms for the states we are interested in, but with two key differences. We are learning terms for both the first and second states in this derivation and thus applying PEM to both of these states as well. We also modify Eq. 13 to train on only data for the first two states, as these are the only states typically observable in experimental data. The initial parameters for this system are given in Table 3, and the formal description of the PEM-UDE system is given by
@@
$\dot{r}$ $=\frac{\Delta}{\pi}+2rv-(\alpha+g_{\text{syn}}s)r+f_{1}(\overset{\rightharpoonup}{u})$ (23a)
$\dot{v}$ $=v(v-\alpha)-w+\bar{\eta}+I_{\text{ext}}+g_{\text{syn}}s(e_{r}-v)-(\pi r)^{2}+f_{2}(\overset{\rightharpoonup}{u})$ (23b)
@@
$f_{1}(\overset{\rightharpoonup}{u})$ $=\phi_{1}s+\phi_{2}rs+\phi_{3}vs+c_{1}$ (23e)
$f_{2}(\overset{\rightharpoonup}{u})$ $=\phi_{4}r+\phi_{5}rv+\phi_{6}rs+\phi_{7}s^{2}+c_{2}$ (23f)
We note, however, that this form is fit only to the sparsity at 50% connectivity (see Fig. 11 in the Supplement) and requires additional tuning to different sparsities. Details on the tuning algorithms are discussed in the Supplemental Information.
@@
$b$ -0.0062 Dimensionless Izhikevich recovery current parameter
@@
Table 3: Parameters of the Izhikevich neuron population and NGNMM, from [chen2022].
@@
$\phi_{1}$ -0.50 Reduction of synaptic activity contribution to firing rate
$\phi_{2}$ -0.12 Rescaling of $g_{\text{syn}}$
$\phi_{3}$ 0.88 Voltage-dependent synaptic activity
$\phi_{4}$ -5.01 Additional scaling term for firing rate contribution to voltage
$\phi_{5}$ 9.80 Non-linear interaction between firing rate and voltage
$\phi_{6}$ -0.05 Non-linear interaction between firing rate and synaptic activity
$\phi_{7}$ -0.27 Quadratic contribution of synaptic dynamics to voltage
$c_{1}$ 0.12 Rescaling of $\Delta$
$c_{2}$ 0.06 Rescaling of $\bar{\eta}$
Table 4: Parameters of the symbolic PEM-UDE learned NGNMM at $p_{c}=0.5$.